OpenStudy (anonymous):

1 2/3 divided by 3/4

OpenStudy (texaschic101):

First lets turn the mixed number into an improper fraction 1 2/3 = 5/3 Now we have : (5/3) / (3/4) when dividing fractions, you " flip " what your dividing by, then multiply 5/3 * 4/3 and just multiply giving us : 20/9 or in mixed form, it is 2 2/9 Just make sure that you " flip " what you are dividing by, and not the other number (5/3) / (3/4) = 5/3 * 4/3 = 20/9 ---correct (5/3) / (3/4) = 3/5 * 3/4 = 9/20 --- wrong any questions ?

OpenStudy (anonymous):

its 6/3 not 5/3

OpenStudy (texaschic101):

1 2/3 = 5/3